# FedOC: Optimizing Global Prototypes with Orthogonality Constraints for Enhancing Embeddings Separation in Heterogeneous Federated Learning

### v1

@@ 3.3 The Suitability of OC for CE Loss
where $\theta_{r^{y},w^{y}}$ represents the angle between $r^{y}$ and $w^{y}$. Thus, $w^{y}$ can be considered as the global prototype $\hat{P}^{y}$ for class $y$. It can be observed that when $\theta_{r^{y},w^{y}}$ approaches 0, the dot product is maximized, thus aligning two vectors in feature space. This alignment imposes an angular constraint on their relationship, which is why CE loss is said to have an angular property. However, CE lacks an explicit mechanism to enforce inter-class separation. This limitation facilitates the overlap of different classes in the feature space, leading to suboptimal discrimination.
Recognizing this limitation, we propose to enhance the discriminative power of CE loss by introducing orthogonality constraints (OC). By enforcing orthogonality between class prototypes, we explicitly enhance inter-class separation, ensuring that the feature representations of different classes remain directionally independent. This approach strengthens the inherent angular properties of CE loss, facilitating enhanced intra-class compactness and inter-class discriminability within each mini-batch.
Figure 3: This image illustrates the FedOC framework involving two clients. From ① to ④: We input the samples, get the representations, then input them into the classifier for prediction, and finally calculate the CE loss. ⑤: The dissimilarity between the feature representation and its corresponding prototype is calculated as a regularization term. ⑥: After the local model is updated, we collect the client prototypes. ⑦: Orthogonality constraints are applied to the global prototypes on the server to enhance angular separation.
@@ 3.4 Local Model Update
In FedOC, the client updates its local model to generate embeddings that are consistent with the global prototypes shared across clients. To maintain directional alignment with the global prototype, a regularization term is added to the local loss function, which encourages feature representations from $k$-th client $r_{k}^{c}$ to align with their respective global prototypes $\hat{P}^{c}$ while minimizing classification error. Then, the client prototype $P_{k}^{c}$ has a more separate margin. Specifically, the loss function is formulated as follows:
$\begin{split}\mathcal{L}_{k}:=&\mathbb{E}_{(x,y)\sim B}\mathcal{L}_{CE}(h_{k}(f_{k}(x;\phi_{k});\theta_{k}),y)+\\ &\lambda_{c}\cdot\mathbb{E}_{(x,y)\sim B}\mathcal{L}_{R}(r_{k}^{y},\hat{P}^{y}),\end{split}$ (11)
where $\lambda_{c}$ is a hyperparameter and the regularization is defined as:
$\mathbb{E}_{(x,y)\sim B}\mathcal{L}_{R}(r_{k}^{y},\hat{P}^{y})=1-\frac{1}{|B|}\sum_{(x,y)\in B}\langle r^{y},\hat{P}^{y}\rangle,$ (12)
where $r^{y}$ represents embedding of the sample. This regularization facilitates the alignment between feature representations and global prototypes across clients, enhancing the global consistency and robustness of the federated model in heterogeneous environments.
3.5 FedOC Framework
Algorithm 1 The learning process of FedOC
@@ 3.5 FedOC Framework
17: Collect local prototype $\{P_{k}\}$ by Eq. 4
18: Send updated prototype $\{P_{k}\}$ to the server
We present the complete workflow of FedOC in Alg. 1 and give an illustration in Fig. 3. The well-trained, class-separable global prototypes are then distributed to clients in the subsequent round to guide client training for improving separability among feature representations.
@@ 4 Convergence Analysis
To analyse the convergence of FedOC, we first introduce some additional notes as in the existing framework. Unless otherwise stated, we always write the client-side loss function $\mathcal{L}_{k}(\mathcal{D}_{k},\omega_{k},\mathcal{P})$ as $\mathcal{L}_{k}$. Since all client-side loss function share the same property, we omit the subscript $k$ and denote the loss function as $\mathcal{L}$. In order to analyse the convergence of FedOC, we first introduce some additional notes as in the existing framework.
@@
$\|\nabla\mathcal{L}_{k}^{t_{1}}-\nabla\mathcal{L}_{k}^{t_{2}}\|_{2}\leq L_{1}\|\omega_{k}^{t_{1}}-\omega_{k}^{t_{2}}\|_{2},$ (13)
where $\forall t_{1},t_{2}>0,k\in\{1,\dots,m\}$. That also means that local objective function is $L_{1}$-Lipschitz smooth. The function can be locally approximated by a quadratic function:
$\mathcal{L}_{k}^{t_{1}}-\mathcal{L}_{k}^{t_{2}}\leq\langle\nabla\mathcal{L}_{k}^{t_{2}},(\omega_{k}^{t_{1}}-\omega_{k}^{t_{2}})\rangle+\frac{L_{1}}{2}\|\omega_{k}^{t_{1}}-\omega_{k}^{t_{2}}\|_{2}^{2},$ (14)
@@
$\mathbb{E}_{\xi_{i}\thicksim D_{i}}[g_{k}^{t}]=\nabla\mathcal{L}(\omega_{k}^{t})=\nabla\mathcal{L}^{t},\forall k\in\{1,2,\ldots,m\},$ (15)
@@
$\mathbb{E}[\left\|g_{k}^{t}-\nabla\mathcal{L}(\omega_{k}^{t})\right\|_{2}^{2}]\leq\sigma^{2},\forall k\in\{1,2,\ldots,m\}.$ (16)
Assumption 3. (Bounded Expectation of Euclidean norm of Stochastic Gradients of the Global Prototype). The expectation of euclidean norm of the stochastic gradient of the global prototype is bounded by $G$:
$\mathbb{E}[\|\nabla\mathcal{L}_{server}\|_{2}]\leq G$
We denote $e\in\{0,1,2,\ldots,E\}$ as the local iteration, and t as the global round. Here, $tE$ indicates the time step before prototype aggregation at $t$ round, and $tE+0$ represents the interval between global prototype training and the first iteration of the $(t+1)$ round. Under above assumptions, we present the theoretical results at the non-convex condition. Since FedOC makes no change to the local model training process, Lemma 1 derived by FedProto [33] still holds.
Lemma 1. Based on Assumption 1 and 2, after clients training one round, the loss of an arbitrary client’s local model is bounded by:
@@
Lemma 2. Based on Assumption 3, after clients receiving the latest global prototype, the loss of an arbitrary client’s local model is bounded by:
$\mathbb{E}\left[\mathcal{L}^{(t+1)E+0}\right]\leqslant\mathbb{E}\left[\mathcal{L}^{(t+1)E}\right]+\lambda\eta G.$ (19)
Theorem 1. Leveraging Lemma 1 and Lemma 2, we derive Theorem 1, which establishes an upper bound on the expected reduction in loss achieved during a single round of local training:
@@
$\frac{\eta^{2}L_{1}E\sigma^{2}}{2}+\lambda\eta G.$
Theorem 2. We extend the analysis to multiple rounds, providing a convergence rate for FedOC in a non-convex setting. For an arbitrary client and any $\epsilon>0$, the following inequality holds:
@@
$\frac{2\lambda G}{2-L_{1}\eta}\leqslant\epsilon$
when $\eta<\frac{2(\epsilon-\lambda G)}{L_{1}\left(\epsilon+E\sigma^{2}\right)}$ and $\lambda<\frac{\epsilon}{G}$. It shows that for an arbitrary client, the model can converge to a solution with a rate that is proportional to $O(1/T)$, where $T$ is the number of communication rounds.
@@ 5 Experiments
Table 1: The test accuracy (%) on three datasets in the pathological and practical settings using the HtFE8 model group.
@@
Datasets Cifar10 Cifar100 Flowers102 Cifar10 Cifar100 Flowers102
FML 82.56 ± 0.14 50.27 ± 0.09 52.17 ± 0.98 91.85 ± 0.24 46.45 ± 0.17 48.51 ± 0.57
LG-FedAvg 83.63 ± 0.09 55.51 ± 0.10 58.90 ± 0.22 92.34 ± 0.06 49.82 ± 0.16 53.65 ± 0.46
FedGen 83.63 ± 0.24 55.37 ± 0.35 59.39 ± 0.19 92.43 ± 0.11 49.90 ± 0.21 54.59 ± 0.34
FedProto 80.63 ± 0.05 51.89 ± 0.31 54.48 ± 0.20 79.32 ± 0.22 42.77 ± 0.13 26.66 ± 0.57
FedKD 83.41 ± 0.66 53.01 ± 1.48 52.61 ± 0.74 91.85 ± 0.31 48.95 ± 1.24 51.44 ± 0.75
FedGH 83.49 ± 0.29 55.25 ± 0.12 60.15 ± 1.08 92.66 ± 0.10 49.52 ± 0.05 54.24 ± 0.50
FedTGP 81.60 ± 0.30 52.17 ± 0.74 52.43 ± 0.64 82.48 ± 1.03 47.96 ± 0.04 44.56 ± 0.45
FedOC 85.71 ± 0.10 61.12 ± 0.15 64.15 ± 1.66 94.01 ± 0.03 55.33 ± 0.23 60.34 ± 0.48
@@ 5.1 Setup
Datasets. To evaluate our model, we conduct experiments on three widely used image classification datasets: CIFAR-10, CIFAR-100 [11] and Flowers102 [25].
Baselines. To evaluate our proposed FedOC, we compare it with seven popular methods that are applicable in HtFL, including LG-FedAvg [19], FML [29], FedGen [44], FedKD [39], FedProto [33], FedGH [41] and FedTGP [43].
@@
Statistical heterogeneity. We conduct extensive experiments with two widely used statistically heterogeneous settings, the pathological setting and the practical setting. For the pathological setting, we assign a fixed number of classes to the client. For the practical setting, we leverage Dirichlet distribution to simulate more realistic class imbalance across clients. The hyperparameter $\alpha$ controls the strength of heterogeneity. Notably, a smaller $\alpha$ implies a higher nonIID in the data distribution among clients.
Training configuration. Our experiments are conducted on an x86_64 architecture with Ubuntu as the operating system. The platform is equipped with 8 NVIDIA V100 GPUs, each with 32 GB of memory, and CUDA version 12.2. Unless explicitly specified, we use the following settings. The federated learning setup includes a default configuration of 20 clients and the client participation ratio $\rho$ = 1. Both the client and server employ the SGD optimizer with a learning rate of 0.01, and both undergo training for a single epoch. The batch size is set to 32 for all experiments. For model heterogeneity robustness, we use the HMG8 by default. For statistical heterogeneity testing, under the pathological setting, we distribute unbalanced data of 2/10/10 classes to each client from a total of 10/100/102 classes on Cifar10/Cifar100/Flowers102 datasets. In the practical setting, a Dirichlet distribution with $\alpha$ = 0.05 is used to simulate real life scenario. Each client’s private data is divided into a training set (75%) and a test set (25%). Each algorithm is trained three times, with each training run consisting of 100 epochs. For the evaluation, we count the best accuracy achieved in each training run, and the final result is calculated as the mean and variance of the accuracy of the three runs.
Evaluation Metrics. We measure the average test accuracy across all clients’ local models. We further evaluate the robustness of our algorithm under varying client participation rates, degrees of non-IID data, and feature dimension $K$.
@@ 5.2 Performance Comparison
The test accuracy of all methods across three datasets is presented in Tab. 1. FedOC consistently outperforms all baselines on these datasets, achieving up to a 15.78% gain over FedTGP in practical settings and 11.72% in pathological settings on Flowers102. This improvement stems from FedOC’s ability to reduce similarity among class representations within the feature space, enhancing classification efficacy more effectively than methods that merely increase Euclidean distance for class separation in high-dimensional spaces. Moreover, FedOC excels in more challenging tasks, where a larger number of classes entails more client prototypes. Its superior performance with greater class diversity is largely due to its orthogonality constraints, which improve inter-class separation within feature spaces.
Table 2: The test accuracy (%) on Cifar100 with three different HMGs to test robustness to heterogeneous models.
@@ 5.3 Impact of Model Heterogeneity
To examine the impact of model heterogeneity in HtFL, we assess the performance of FedOC on three additional HMG setttings. We show results in Tab. 2, our findings indicate that all baselines performe well under the HMG3 condition, as the reduced model heterogeneity significantly lessened its adverse impact on performance, while all methods perform worse with larger model heterogeneity. As the degree of model heterogeneity rises, the dominance of FedOC becomes more pronounced, up to 10.12% at HMG5 and 8.13% at HMG10. In addition, FedOC also has the smallest performance reduction when model heterogeneity increases, which shows that FedOC is robust to model heterogeneity.
@@ 5.4 Robustness to Participation Rate
To compare FedOC against baselines under varying total client numbers $M$ and client participation rates $\rho$, we design three distinct settings. The results, shown in Tab. 3, ensure a consistent number of participating clients per round. FedOC aims to promote directional independence, which is seamlessly integrated with CE loss. This separation minimizes inter-class overlap, thereby enabling the model to achieve robust generalization across clients with varying participation rates.
Table 3: The test accuracy (%) on Cifar100 with three participation rate settings. “-” means the baseline can’t converge.
@@
FedTGP 42.55 ± 0.65 26.02 ± 0.54 27.27 ± 0.23
FedOC 54.88 ± 0.45 44.86 ± 0.11 43.22 ± 0.45
Table 4: Test the accuracy (%) of two data distributions with various degrees of non-IID on the Cifar100 dataset.
@@
5 classes/client 20 classes/client dir=0.1 dir=0.01

### v2

@@ 3.3 The Suitability of Orthogonality for CE Loss
where $\theta_{r^{y},w^{y}}$ represents the angle between $r^{y}$ and $w^{y}$. Thus, when the angle between the feature representation $r^{y}$ and the weight vector $w^{y}$ approaches zero, their inner product is maximized, indicating directional alignment. Meanwhile, since the feature representation $r^{y}$ is explicitly aligned with the global prototype $\hat{P}^{y}$ during training, the three components ($r^{y}$, $w^{y}$, and $\hat{P}^{y}$) continuously align with each other in the feature space. However, the Cross-Entropy (CE) loss inherently lacks an explicit mechanism to enforce angular separation between different classes, leading to potential overlap in feature representations and consequently suboptimal discriminative performance. To overcome this limitation, we propose orthogonality regularization (OR) to explicitly enlarge angular separation among global prototypes. By enforcing orthogonality among global prototypes, OR significantly reduces inter-class overlap, thereby ensuring robust directional consistency among $r^{y}$, $w^{y}$, and $\hat{P}^{y}$, and ultimately improving the discriminative capability of the classifier in heterogeneous federated environments.
Figure 3: This image illustrates the FedORGP framework involving two clients. From ① to ④: We input the samples, get the representations, then input them into the classifier for prediction, and finally calculate the CE loss. ⑤: The dissimilarity between the feature representation and its corresponding prototype is calculated as a regularization term. ⑥: After the local model is updated, we collect the client prototypes. ⑦: Orthogonality regularization are applied to the global prototypes on the server to enhance angular separation.
@@ 3.4 Local Model Update
In FedORGP, the client updates its local model to generate embeddings that are consistent with the global prototypes shared across clients. To maintain directional alignment with the global prototype, a regularization term is added to the local loss function, which encourages feature representations from $k$-th client $r_{k}^{c}$ to align with their respective global prototypes $\hat{P}^{c}$ while minimizing classification error. Then, the client prototype $P_{k}^{c}$ has a more separate margin. Specifically, the loss function is formulated as follows:
$\begin{split}\mathcal{L}_{k}:=&\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\mathcal{L}_{CE}(h_{k}(f_{k}(x;\phi_{k});\theta_{k}),y)+\\ &\lambda_{c}\cdot\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\mathcal{L}_{R}(f_{k}(x;\phi_{k}),\hat{P}^{y}),\end{split}$ (8)
where $\lambda_{c}$ is a hyperparameter. This regularization $\mathcal{L}_{R}$ facilitates the alignment between feature representations and global prototypes across clients, enhancing the global consistency and robustness of the federated model in heterogeneous environments.
3.5 FedORGP Framework
Algorithm 1 The learning process of FedORGP
@@ 3.5 FedORGP Framework
17: Collect local prototypes $\{P^{c}_{k}|c\in\mathcal{C}_{k}\}$ by Eq. 2
18: Send local prototypes $\mathcal{P}_{k}$ to the server
We present the complete workflow of FedORGP in Alg.1 and give an illustration in Fig.3. The well-trained, class-separable global prototypes are then distributed to clients in the subsequent round to guide client training for improving separability among feature representations.
@@ 4 Convergence Analysis
To analyse the convergence of FedORGP, we first introduce some additional notes as in the existing framework. Unless otherwise stated, we always write the client-side loss function $\mathcal{L}_{k}(\mathcal{D}_{k},\omega_{k},\mathcal{P})$ as $\mathcal{L}_{k}$. In order to analyse the convergence of FedORGP, we first introduce some additional assumptions [32, 42, 4].
@@
$\|\nabla\mathcal{L}_{k}^{t_{1}}-\nabla\mathcal{L}_{k}^{t_{2}}\|_{2}\leqslant L_{1}\|\omega_{k}^{t_{1}}-\omega_{k}^{t_{2}}\|_{2},$ (9)
where $\forall t_{1},t_{2}>0,k\in\{1,\dots,m\}$. That also means that local objective function is $L_{1}$-Lipschitz smooth:
$\mathcal{L}_{k}^{t_{1}}-\mathcal{L}_{k}^{t_{2}}\leqslant\langle\nabla\mathcal{L}_{k}^{t_{2}},(\omega_{k}^{t_{1}}-\omega_{k}^{t_{2}})\rangle+\frac{L_{1}}{2}\|\omega_{k}^{t_{1}}-\omega_{k}^{t_{2}}\|_{2}^{2},$ (10)
@@
$\mathbb{E}_{\xi_{i}\thicksim D_{i}}[g_{k}^{t}]=\nabla\mathcal{L}_{k}(\omega_{k}^{t}),\forall k\in\{1,2,\ldots,m\},$ (11)
@@
$\mathbb{E}[\left\|g_{k}^{t}-\nabla\mathcal{L}_{k}(\omega_{k}^{t})\right\|_{2}^{2}]\leqslant\sigma^{2},\forall k\in\{1,2,\ldots,m\}.$ (12)
Assumption 3. (Bounded Gradient). The stochastic gradient of the global prototype is bounded by $G$:
$\mathbb{E}[\|\nabla\mathcal{L}_{server}\|_{2}]\leqslant G$
We denote $e\in\{0,1,2,\ldots,E\}$ as the local iteration, and t as the global round. Here, $tE$ indicates the time step before orthogonality regularization at $t$ round, and $tE+0$ represents the interval between global prototype training and the first iteration of the $(t+1)$ round. Under above assumptions, we present the theoretical results at the non-convex condition. Since all client-side loss function share the same property, we omit the subscript $k$ and denote the loss function as $\mathcal{L}$. The detailed proof process is in the Appendix.
Lemma 1. Based on Assumption 1 and 2, after clients training one round, the local client model loss satisfies [32]:
@@
Lemma 2. Based on Assumption 3, we further analyze how the introduction of global prototype orthogonality regularization influences the client-side loss:
$\mathbb{E}\left[\mathcal{L}^{(t+1)E+0}\right]\leqslant\mathbb{E}\left[\mathcal{L}^{(t+1)E}\right]+\lambda_{c}\eta G.$ (15)
Theorem 1. Leveraging Lemma 1 and Lemma 2, we establish our main theoretical convergence results of the loss reduction over multiple steps in local training:
@@
$\frac{\eta^{2}L_{1}E\sigma^{2}}{2}+\lambda_{c}\eta G.$
Theorem 2. We extend the analysis to multiple rounds, providing a convergence rate for FedORGP in a non-convex setting. For an arbitrary client and any $\epsilon>0$, the following inequality holds:
@@
$\frac{2\lambda_{c}G}{2-L_{1}\eta}\leqslant\epsilon$
when $\eta<\frac{2(\epsilon-\lambda_{c}G)}{L_{1}\left(\epsilon+E\sigma^{2}\right)}$ and $\lambda_{c}<\frac{\epsilon}{G}$. It shows that the model can converge to a solution with a rate that is proportional to $O(1/T)$, where $T$ is the number of communication rounds.
@@ 5 Experiments
Table 1: The test accuracy (%) on four datasets in the pathological and practical settings using the HMG8 model group.
@@
Datasets Cifar-10 Cifar-100 Flowers102 TinyImagenet Cifar-10 Cifar-100 Flowers102 TinyImagenet
FML 82.56 ± 0.14 50.27 ± 0.09 52.17 ± 0.98 33.37 ± 0.34 91.85 ± 0.24 46.45 ± 0.17 48.51 ± 0.57 37.67 ± 0.19
LG-FedAvg 83.63 ± 0.09 55.51 ± 0.10 58.90 ± 0.22 34.99 ± 0.29 92.34 ± 0.06 49.82 ± 0.16 53.65 ± 0.46 38.35 ± 0.27
FedGen 83.63 ± 0.24 55.37 ± 0.35 59.39 ± 0.19 35.18 ± 0.12 92.43 ± 0.11 49.90 ± 0.21 54.59 ± 0.34 38.44 ± 0.14
FedProto 80.63 ± 0.05 51.89 ± 0.31 54.48 ± 0.20 33.78 ± 0.20 79.32 ± 0.22 42.77 ± 0.13 26.66 ± 0.57 24.57 ± 0.08
FedKD 83.41 ± 0.66 53.01 ± 1.48 52.61 ± 0.74 34.70 ± 0.91 91.85 ± 0.31 48.95 ± 1.24 51.44 ± 0.75 39.13 ± 0.36
FedGH 83.49 ± 0.29 55.25 ± 0.12 60.15 ± 1.08 35.04 ± 0.16 92.66 ± 0.10 49.52 ± 0.05 54.24 ± 0.50 38.67 ± 0.11
FedTGP 84.75 ± 0.06 53.12 ± 0.24 58.60 ± 0.14 32.41 ± 0.09 92.26 ± 0.68 49.09 ± 0.16 56.62 ± 0.40 37.14 ± 0.16
FedORGP 85.71 ± 0.10 61.12 ± 0.15 64.15 ± 1.66 37.04 ± 0.56 94.01 ± 0.03 55.33 ± 0.23 60.34 ± 0.48 40.15 ± 0.40
@@ 5.1 Setup
Datasets. To evaluate our model, we conduct experiments on four image classification datasets: CIFAR-10, CIFAR-100 [14], Flowers102 [26] and TinyImagenet [3].
Baselines. To evaluate our proposed FedORGP, we compare it with seven popular methods that are applicable in HtFL, including LG-FedAvg [20], FML [29], FedGen [45], FedKD [39], FedProto [32], FedGH [42] and FedTGP [44].
@@
Statistical heterogeneity. We conduct extensive experiments with two widely used statistically heterogeneous settings, the pathological setting and the practical setting [25, 45, 21]. For the pathological setting, we assign a fixed number of classes to the client. For the practical setting, we leverage Dirichlet distribution to simulate more realistic class imbalance across clients. The hyperparameter $\alpha$ controls the strength of heterogeneity. Notably, a smaller $\alpha$ implies a higher non-IID data distribution among clients.
Training configuration. Our experiments are conducted on an x86_64 architecture with Ubuntu as the operating system. The platform is equipped with 8 NVIDIA V100 GPUs, each with 32 GB of memory, and CUDA version 12.2. Unless explicitly specified, we use the following settings. The federated learning setup includes a default configuration of 20 clients and the client participation ratio $\rho$ = 1. Both the client and server employ the SGD optimizer with a learning rate of 0.01, and both undergo training for a single epoch. The batch sizes for client $B$ and server $B_{p}$ are both set to 32. For model heterogeneity, we use the HMG8 by default. For statistical heterogeneity testing, under the pathological setting, we distribute unbalanced data of 2/10/10/20 classes to each client from a total of 10/100/102/200 classes on Cifar-10/Cifar-100/Flowers102/TinyImagenet datasets. In the practical setting, a Dirichlet distribution with $\alpha$ = 0.05 is used to simulate real life scenario. Each client’s private data is divided into a training set (75%) and a test set (25%). Each algorithm is trained three times, with each training run consisting of 100 epochs. For the evaluation, we count the best accuracy achieved in each training run, and the final result is calculated as the mean and variance of the accuracy of the three runs.
Evaluation Metrics. We measure the average test accuracy (%) across all client models. We further evaluate the robustness of our algorithm under varying client participation rates, degrees of non-IID data, and feature dimension $K$.
@@ 5.2 Performance Comparison
The test accuracy of all methods across four datasets is presented in Tab. 1. FedORGP consistently outperforms all baselines on these datasets, achieving up to a 8% gain over FedTGP in pathological settings and 6.24% in practical settings on Cifar-100. This improvement is due to the fact that FedORGP can separate features in terms of angles, which can be well integrated with the CE loss, thereby improving classification results more effectively than methods which only increase the Euclidean distance to perform prototype separation in high-dimensional space. In addition, FedORGP is also more efficient. While FedTGP requires 100 epochs of training on the server, FedORGP requires only 1 epoch of training with orthogonal regularisation.
Table 2: Test on Cifar-100 (10 classes per client) with three different HMGs to test robustness to heterogeneous models.
@@ 5.3 Impact of Model Heterogeneity
To examine the impact of model heterogeneity in HtFL, we assess the performance of FedORGP on three additional HMG settings. We show results in Tab. 2, our findings indicate that all baselines perform well under the HMG3 condition, as the reduced model heterogeneity significantly lessened its adverse impact on performance, while all methods perform worse with larger model heterogeneity. As the degree of model heterogeneity rises, the advantages of FedORGP becomes more pronounced, up to 10.12% at HMG5 and 8.13% at HMG10. In addition, FedORGP also has the smallest performance reduction when model heterogeneity increases, which shows that FedORGP is robust to model heterogeneity.
@@ 5.4 Robustness to Participation Rate
To compare FedORGP against baselines under varying total client numbers $M$ and client participation rates $\rho$, we design three distinct settings. The results, shown in Tab. 3, ensure a consistent number of participating clients per round. FedORGP aims to promote directional independence, which is seamlessly integrated with CE loss. This separation minimizes inter-class overlap, thereby enabling the model to achieve robust generalization across clients with varying participation rates.
Table 3: Test on Cifar-100 (10 classes per client) under three participation rate settings with HMG8. “-” means the baseline can’t converge.
@@
FedTGP 44.87 ± 0.58 37.51 ± 0.36 36.75 ± 0.42
FedORGP 54.88 ± 0.45 44.86 ± 0.11 43.22 ± 0.45
Table 4: Test on two data distributions under various degrees of non-IID on the Cifar-100 dataset with HMG8.
@@
5 classes/client 20 classes/client $\alpha$=0.1 $\alpha$=0.01